Today we found the green beam from the end was totally missing at the vertex.
- What we found was very weak green beam at the end. Unhappy.
- We removed the PBS. We should obtain the beam for the fiber from the rejection of the (sort of) dichroic separator although the given space is not large.
- The temperature controller was off. We turned it on again.
- We found everything was still misaligned. Aligned the crystal, aligned the Faraday for the green.
- Aligned the last two steering mirrors such that we hit the approximate center of the ETMX and the center of the ITMX.
- Made the fine alignment to have the green beam at the PSL table.
The green beam emerged from the chamber looks not so round as there is a clipping at an in-vac steering.
We will make the thorough realignment before closing the tank.